how do you simplify 15.35 into a fraction?
Answer by stanbon(75887)

how do you simplify 15.35 into a fraction?
----
15.35 = 15.35/1
-----
Multiply numerator and denominator by 100 to get:
= 1535/100
= 307/20
============
Cheers,
Stan H.
============